March 8, 2023
Return to the calendar.
RC Meeting Information
Current RC: Iuliia Skorodumina.
No RC meeting today.
Tuesday Day (03/07). The beam was taken away at 10 am for the 4th pass separator phasing to continue. No beam for the rest of the shift. Shortly after the beam was gone, there were HMS Q3 interlock and HMS dipole interlock, which were resolved by the shift crew (https://logbooks.jlab.org/entry/4144408 and https://logbooks.jlab.org/entry/4144414). Day Shift summary: https://logbooks.jlab.org/entry/4144353 .
Tuesday Swing (03/07). The 4th pass separator has been phased (https://logbooks.jlab.org/entry/4144489) and MCC started sending us some beam at around 4:30pm. It was decided to check the harp scans and the hole position. The harp scans looked fine (https://logbooks.jlab.org/entry/4144547 and https://logbooks.jlab.org/entry/4144548) and the hole was centered (https://logbooks.jlab.org/entry/4144583). When switching back to production, there was an issue with the raster GUI, which followed by SHMS PreScaledTrig2 alarm (https://logbooks.jlab.org/entry/4144597). For the latter, the alarm limits were adjusted (https://logbooks.jlab.org/entry/4144600). Data taking resumed, but the beam was not stable (https://logbooks.jlab.org/entry/4144659 and https://logbooks.jlab.org/entry/4144750). Swing shift summary: https://logbooks.jlab.org/entry/4144513 .
Wednesday Owl (03/08). Taking production data at 50 uA. Fast Feedback system was turned back on at 3:40am (https://logbooks.jlab.org/entry/4144942). We have reached (and exceeded) the planned integrated charge for the kinematics with Pm=800 MeV/c (we have collected a total of 16172 mC of the planned 15696 mC). Owl Shift summary: https://logbooks.jlab.org/entry/4144804 .
Wednesday Day (03/08). Shift started with the configuration change. We rotated HMS to 61.3 deg and changed HMS momentum to +2.0474 GeV. We are now in the setting with Pm=900 MeV/c, and the plan is to stay there for the rest of the experiment and to collect 18720 mC of total integrated charge. When changing the HMS momentum, HMS Q2 stopped responding (https://logbooks.jlab.org/entry/4145080) and we called Amy Comer. She managed to restore the communication with the magnet remotely (https://logbooks.jlab.org/entry/4145093), and we started taking production data at this new kinematic setting.
Current Run Plan
We are now in the setting with Pm=900 MeV/c, and the plan is to stay there for the rest of the experiment and to collect 18720 mC of the total integrated charge.